OpenStudy (anonymous):

A tank is is half full of oil that has a density of 900 kg/m3. Find the work W required to pump the oil out of the spout. (Use 9.8 m/s2 for g. Assume r = 3 m and h = 1 m.)

OpenStudy (aum):

Find the volume of oil. Multiply by density to find its mass. Multiply by g to find its weight. Have you done calculus yet? If not we have to use some another method to find the work done.

OpenStudy (anonymous):

Work = Force x Distance, right? I did found the force to equal 158760pi, but I have no idea how to find the distance in this case.

OpenStudy (aum):

Have you been taught calculus yet?

OpenStudy (anonymous):

Yes.

OpenStudy (aum):

Is h = 1m the height of the oil tank or height of the oil in the tank?

OpenStudy (anonymous):

http://www.webassign.net/scalcet7/6-4-022-alt.gif This is the picture of the tank.

OpenStudy (aum):



OpenStudy (aum):

\[ x^2 + y^2 = 3^2 = 9\\ x^2 = 9 - y^2 \\ dV = \pi x^2 dy \\ dF = \pi x^2 dy * 900 * 9.8 ~N = 27,708x^2dy = 27,708(9-y^2)dy\\ dW = F * d = 27,708(9-y^2)dy * (3-y+1) = 27,708(9-y^2)(4-y)dy\\ \]

OpenStudy (aum):

Work done = \[\large \int_{-3}^027,708(9-y^2)(4-y)dy \]

OpenStudy (anonymous):

Can you explain to me why it's -3 to 0 as opposed to 0 to 3?

OpenStudy (aum):

The tank is half full. The origin is at the center of the sphere with positive x axis to the right and positive y-axis up.

OpenStudy (anonymous):

That makes sense! Thank you for your help!

OpenStudy (aum):

You are welcome.

OpenStudy (aum):

Do you have the answer to this?

OpenStudy (anonymous):

Yes! 2556063 J